OpenStudy (anonymous):

Find a polynomial (of degree 3) with the given zeros. 1, 0, -1

myininaya (myininaya):

f(x)=(x-1)(x-0)(x-(-1)) f(x)=(x-1)x(x+1) f(x)=x(x-1)(x+1) f(x)=x(x^2-1)